OpenStudy (anonymous):

On segment AB, C is between A and B. Given AC = 12 and CB = 16, which is proportional to the ratio of (BC − AC) to AB?

OpenStudy (anonymous):

16 - 12 = 4 16 + 12 = 28 4 / 28 = 1/7

OpenStudy (anonymous):

Thankyou!